OpenStudy (anonymous):

there are 20 dogs and 30 cats. All dogs weigh over 15 pounds. Only 1/6 of all the cats weighs over 15 pounds. One of the animals is chosen at random, and it weights over 15 pounds. What is the probability that the animal chosen is a dog?

OpenStudy (anonymous):

30*(1/6)+20 5+20 25 one out of 25

OpenStudy (anonymous):

The options are 2/5, 4/5, 3/8, and 1/4

OpenStudy (anonymous):

=/

OpenStudy (anonymous):

Although your explanation does make sense... do you mean 2/5?

OpenStudy (ranga):

Number of dogs = 20 Total number of animals = Total dogs + Total cats = 20 + 30 = 50 If an animal is chosen at random, the probability that it is a dog = 20 / 50 = 2/5 All the other information about weight is extraneous information intended to confuse student. The basic question is if an animal is chosen at random what is the probability that it is a dog and the answer is 2/5.

OpenStudy (anonymous):

Ughh I hate when they do that lol thank you though, I understand it now.

OpenStudy (ranga):

you are welcome!